OpenStudy (anonymous):

simply square root 294x to the second power

OpenStudy (anonymous):

Is the question this? \[\sqrt{294x^2}\] Then, 294=2*147=2*7*21=2*3*7*7 \[\sqrt{294x^2}=7x\sqrt{6}\]

OpenStudy (anonymous):

yes it is. and thanks

OpenStudy (anonymous):

Welcome! ^^